If A and B are events such that P(A) = 0.6, P(B|A) = 0.3 and P(A ∪ B) = 0.72. Are A and B independent, mutually exclusive or both?


1
Expert's answer
2021-03-01T06:41:07-0500

Solution:


We are given "P(A), P(B|A), P(A+B)"

If "P(B|A)=P(B)" We can say that events are independen.Therefore, we need to find "P(B)". To do so we need to remember the formula of sum of events:


"P(A+B) = P(A) + P(B) - P(A*B)"

"P(B) = P(A+B) + P(A*B) - P(A)"


And "P(A*B) = P(A)*P(B|A)= 0.6 * 0.3 = 0.18", so:


"P(B) = 0.72 + 0.18 - 0.6=0.3"

"P(B) = 0.3 = P(B|A)," therefore the events are independent.


For events to be mutually exclusive, they can not occur at the same time, so following equation has to occur:


"P(A*B) = 0",


but we previously found: "P(A*B) = 0.18", therefore the events are not mutually exclusive.


Answer:


Events are independent, but not mutually exclusive